# Towards Open-ended Visual Quality Comparison

### 26 Feb 2024 v1

@@ Abstract
Comparative settings (e.g. pairwise choice, listwise ranking) have been adopted by a wide range of subjective studies for image quality assessment (IQA), as it inherently standardizes the evaluation criteria across different observers and offer more clear-cut responses. In this work, we extend the edge of emerging large multi-modality models (LMMs) to further advance visual quality comparison into open-ended settings, that 1) can respond to open-range questions on quality comparison; 2) can provide detailed reasonings beyond direct answers. To this end, we propose the Co-Instruct. To train this first-of-its-kind open-source open-ended visual quality comparer, we collect the Co-Instruct-562K dataset, from two sources: (a) LMM-merged single image quality description, (b) GPT-4V “teacher” responses on unlabeled data. Furthermore, to better evaluate this setting, we propose the MICBench, the first benchmark on multi-image comparison for LMMs. We demonstrate that Co-Instruct not only achieves 30% higher superior accuracy than state-of-the-art open-source LMMs, but also outperforms GPT-4V (its teacher), on both existing related benchmarks and the proposed MICBench.
@@ 1 Introduction
While comparison has widely-recognized significance for IQA, existing related datasets [35] and methods [64, 28] are generally based on overall quality comparison and have not extended to the open-ended scenarios; on the other hand, open-source LMMs [51, 25, 4] are usually only fine-tuned with single image instruction tuning datasets [26, 13, 1] and proved to lack enough capability even on two image comparison settings [68, 70]. While these gaps have clearly indicated the need of a specific instruction tuning dataset for visual quality comparison, it is too expensive to collect such a dataset from human. To avoid costly human labors, we propose an alternative strategy to collect the training dataset, named Collaborative Instruction Tuning from Weak Supervisors (Co-Instruct). Specifically, we adopt two non-perfect supervisors: 1) Merge2Compare. Originated from single image human quality descriptions of 19K images in Q-Pathway [51], we randomly match them into 100K groups with removing the most similar descriptions with an text embedding model [44]. Then, similar as the construction of LLaVA-150K [26], we prompt a single-modal LLM [16] to compare multiple human descriptions in a group, and “merge” them into 100K pseudo comparisons (as illustrated in Fig. 3(a)). 2) Teach2Compare. Observing that GPT-4V has especially high accuracy on pairwise settings [70, 68] among existing LMMs, following existing practices [2, 17], we leverage GPT-4V responses to further expand our dataset. We collect 9K unlabeled images and randomly match into 30K image groups (2-4 images per group), and obtain GPT-4V responses on both caption-like general comparisons and question-answer pairs for comparisons. By integrating Q-Instruct-200K [51] (on single images), Merge2Compare, and Teach2Compare, we construct the Co-Instruct-562K, the first instruction tuning dataset designed for open-ended multi-image quality comparison.
@@
As we need to feed multiple images together during instruction tuning, adopting the most popular LLaVA [26, 39, 25] structure that linearly projects visual embeddings will exceed the context window of the language models [42] and cause errors. Henceforth, we adopt an alternative visual abstractor structure [57] to first reduce visual token length (from 1025 to 65 tokens per image), and then concatenate them with text embeddings to pass to language decoders. By learning from the Co-Instruct-562K dataset and the specially-designed input structure, we present the Co-Instruct, with up to 86% improvements than its baseline [57], and 61% better than the state-of-the-art open-source LMM. More importantly, despite using GPT-4V as one of its teachers, it still surpasses the GPT-4V teacher in a variety of multi-choice question (MCQ) benchmarks, and matches GPT-4V ability in scenarios requiring detailed language reasonings.
@@ 2.2 LMMs for Visual Quality Evaluation
| # Total Images | 19K (shared, both using Q-Pathway images) | 9K (using shared images) | 28K | ||
| # Total Data Items | 202K | 100K | 30K | 230K | 562K |
@@ 4 The Co-Instruct Model
Most state-of-the-art LMMs [25, 39, 5] have adopted the simple projector that keeps a large number of tokens (e.g. 1025). This structure is not friendly to multi-image scenarios: passing only two images will exceed the max length of LLaVA [26] (2048), and four images will exceed the context window of LLaMA-2 [42] (4096). Thus, we adopt another widely-used abstractor [4, 56, 57] structure to reduce token numbers before feeding the visual embeddings to LLM, so as to easily adapt to multi-image scenarios.
@@
Typical single-image instruction tuning usually does not care about “position of images”. Most approaches [69, 25, 4] directly pile all images before texts (<img>(<img>)<text>). While this will not cause much troubles if we query about joint information from multiple images, it is observed to introduce hallucinations across images during comparison. To solve this, we propose an image-text interleaved format for multi-image training, that each image is started with explicit text to identify its nominal:
@@ 6.1 Implementation Details
The Co-Instruct is fine-tuned after the released checkpoint of mPLUG-Owl2 [57], with LLaMA-2 [42] as LLM and CLIP-ViT-L14 [36] as visual embedding module. Images are padded to square and then resized to before fed into the model. The learning rate is set as , with two epochs under batch size . The final checkpoint is used for evaluation. To avoid over-fitting, only dev subsets of evaluation datasets are used to choose best training hyper-parameters, where the final reported results are from non-overlapped test subsets. All parameters are updated during training, costing in total 25 hours on 8*NVIDIA A100 GPUs.
@@ 6.3 Results on Existing Evaluation Settings
Q-Bench-A2 is a benchmark setting for general and detailed visual quality comparison with detailed reasonings on image pairs. Consisting of 499 image pairs, it employs GPT to evaluate LMM responses against the golden expert-labeled comparisons on Completeness, Precision, and Relevance. As listed in Tab. 3, the Co-Instruct can remarkably improve the Completeness (+57%) and Precision (+59%) of the comparative outputs than the w/o comparative data version, but still falls a little bit behind GPT-4V on the quantitative metrics. This might be because outputs from GPT-4V are more than twice as long as Co-Instruct outputs, while GPT evaluation is observed [22] to be in favor of longer text outputs. To validate where this potential bias occurs here, we qualitatively visualize the result of different LMMs in Fig. 8. As shown in the figure, the baseline open-source LMM even confuses the information from the two images, and Gemini-Pro makes rather poor detailed reasonings. For GPT-4V, with the longest outputs among all LMMs, even it provides an incorrect overall judgment, it still gets the highest precision score (on par with Co-Instruct which has no obvious mistakes), validating the aforementioned bias towards longer outputs.
@@ 6.4 Results on MICBench
| (Feb/2024) InternLM-XComposer2-VL (interpolate to 5120) | 62.95% | 47.29% | 52.02% | 55.70% | 46.51%* | 51.76% |
@@
As is shown in Tab. 6, Co-Instruct provides very competitive accuracy on open-question quality comparison among three/four images, 5.7% better than GPT-4V (best existing) and 6.4% more accurate than non-expert human; open-source LMMs even struggle to obtain 50% accuracy on this setting. Specifically, for InternLM-XComposer2-VL, without reduced tokens, it cannot directly run on MICBench-MCQ, as all groups of four exceeds its original length. To evaluate it, we interpolate its context length from 4096 to 5120. Nonetheless, we still notice a significant performance degradation on its groups-of-four answers (on interpolated positions, labeled with *), demonstrating the limitations of long visual tokens on multi-image capability. This phenomenon highlights the importance to reduce visual tokens (Sec. 4) to adapt to multi-image scenarios.

### 04 Mar 2024 v2

@@ Abstract
Comparative settings (e.g. pairwise choice, listwise ranking) have been adopted by a wide range of subjective studies for image quality assessment (IQA), as it inherently standardizes the evaluation criteria across different observers and offer more clear-cut responses. In this work, we extend the edge of emerging large multi-modality models (LMMs) to further advance visual quality comparison into open-ended settings, that 1) can respond to open-range questions on quality comparison; 2) can provide detailed reasonings beyond direct answers. To this end, we propose the Co-Instruct. To train this first-of-its-kind open-source open-ended visual quality comparer, we collect the Co-Instruct-562K dataset, from two sources: (a) LLM-merged single image quality description, (b) GPT-4V “teacher” responses on unlabeled data. Furthermore, to better evaluate this setting, we propose the MICBench, the first benchmark on multi-image comparison for LMMs. We demonstrate that Co-Instruct not only achieves in average 30% higher accuracy than state-of-the-art open-source LMMs, but also outperforms GPT-4V (its teacher), on both existing related benchmarks and the proposed MICBench.
@@ 1 Introduction
While comparison has widely-recognized significance for IQA, existing related datasets [35] and methods [63, 28] are generally based on overall quality comparison and have not extended to the open-ended scenarios; on the other hand, open-source LMMs [50, 25, 4] are usually only fine-tuned with single image instruction tuning datasets [26, 13, 1] and proved to lack enough capability even on two image comparison settings [67, 69]. While these gaps have clearly indicated the need of a specific instruction tuning dataset for visual quality comparison, it is too expensive to collect such a dataset from human. To avoid costly human labors, we propose an alternative strategy to collect the training dataset, named Collaborative Instruction Tuning from Weak Supervisors (Co-Instruct). Specifically, we adopt two non-perfect supervisors: 1) Merge2Compare (Fig. 3(a)). Originated from single image quality descriptions on 19K images as labeled by human in the Q-Pathway [50] dataset, we randomly match them into 100K groups (2-4 images per group) with removing the most similar descriptions with an text embedding model [44]. Then, we prompt a single-modal large language model (LLM) [16] to compare multiple human descriptions in a group, and “merge” them into 100K pseudo comparisons. 2) Teach2Compare (Fig. 3(b)). Observing that GPT-4V has especially high accuracy on pairwise settings [69, 67] among existing LMMs, following existing practices [2, 17], we leverage GPT-4V responses to further expand our dataset. We collect 9K unlabeled images and randomly match into 30K image groups (also 2-4 images per group), and obtain GPT-4V responses on both caption-like general comparisons and question-answer pairs for comparisons. By integrating Q-Instruct-200K [50] (on single images), Merge2Compare, and Teach2Compare, we construct the Co-Instruct-562K, the first instruction tuning dataset for open-ended visual quality comparison.
@@
Moreover, as we need to feed multiple images together to the LLM decoder, adopting the most popular LLaVA [26, 39, 25] structure that linearly projects visual embeddings will exceed the context window of the language models [42] and cause errors. Henceforth, we adopt an alternative visual abstractor structure [56] to first reduce visual token length (from 1,025 to 65 tokens per image), and then concatenate them with text embeddings to pass to language decoders. By learning from the Co-Instruct-562K dataset and the specially-designed input structure, we present the Co-Instruct, with up to 86% improvements than its baseline [56], and 61% better than the state-of-the-art open-source LMM. More importantly, despite using GPT-4V as one of its teachers, it still surpasses the GPT-4V teacher in a variety of multi-choice question (MCQ) benchmarks, and matches GPT-4V ability in scenarios requiring detailed language reasonings.
@@ 2.2 LMMs for Visual Quality Evaluation
[…]
@@ 4 The Co-Instruct Model
Most state-of-the-art LMMs [25, 39, 5] have adopted the simple projector that keeps a large number of tokens (e.g. 1,025). This structure is not friendly to multi-image scenarios: passing only two images will exceed the max length of LLaVA [26] (2,048), and four images will exceed the context window of LLaMA-2 [42] (4,096). Thus, we adopt another widely-used abstractor [4, 55, 56] structure to reduce token numbers before feeding the visual embeddings to LLM, so as to easily adapt to multi-image scenarios.
@@
Typical single-image instruction tuning usually does not care about “position of images”. Most approaches [68, 25, 4] directly pile all images before texts (<img>(<img>)<text>). Under this piling, multiple images are not separates and LMMs might confuse the information from different images and fail to compare well (see baseline result in Fig. 8). To solve this, we propose an image-text interleaved format for multi-image training, that each image is started with explicit text to identify its nominal:
@@ 6.1 Implementation Details
The Co-Instruct is fine-tuned after the released checkpoint of mPLUG-Owl2 [56], with LLaMA-2 [42] as LLM and CLIP-ViT-L14 [36] as visual embedding module. Images are padded to square and then resized to before fed into the model. The learning rate is set as -, with two epochs under batch size . The final checkpoint is used for evaluation. To avoid over-fitting, only dev subsets of evaluation datasets are used to choose best training hyper-parameters, where the final reported results are from non-overlapped test subsets. All parameters are updated during training, costing in total 25 hours on 8*NVIDIA A100 GPUs.
@@ 6.3 Results on Existing Evaluation Settings
Q-Bench-A2 is a benchmark setting for general and detailed visual quality comparison with detailed reasonings on image pairs. Consisting of 499 image pairs, it employs GPT to evaluate LMM responses against the golden expert-labeled comparisons on Completeness, Precision, and Relevance. As listed in Tab. 3, the Co-Instruct can remarkably improve the Completeness (+57%) and Precision (+59%) of the comparative outputs than the w/o comparative data version, but still falls a little bit behind GPT-4V on the quantitative metrics. This might be because outputs from GPT-4V are more than twice as long as Co-Instruct outputs, while GPT evaluation used here is observed [22] to be in favor of longer text outputs. To further analyze this potential bias, we qualitatively visualize the result of different LMMs in Fig. 8. As shown in the figure, the baseline open-source LMM even confuses the information from the two images, and Gemini-Pro makes rather poor detailed reasonings. For GPT-4V, it generates the longest outputs among all LMMs, which might be the reason that it gets a relatively high precision score even its outputs are not totally correct. In short, the capability of Co-Instruct in reasoning-related comparisons can match that of GPT-4V, while significantly surpassing other existing LMMs.
@@ 6.4 Results on MICBench
| (Feb/2024) InternLM-XComposer2-VL (length: 40965120) | 62.95% | 47.29% | 52.02% | 55.70% | 46.51%* | 51.76% |
@@
As is shown in Tab. 6, Co-Instruct provides very competitive accuracy on open-question quality comparison among three/four images, 5.7% better than GPT-4V (best existing) and 6.4% more accurate than non-expert human; open-source LMMs even struggle to obtain 50% accuracy on this setting. It is also noteworthy that LLaVA series and XComposer2-VL’s original context lengths are not enough for four images as they have not reduced visual token numbers, so we have to extend their context windows to evaluate them for MICBench. Consequentially, all these models have experienced notably worse accuracy on groups of four (on extended context length) than groups of three (within original context length), as noted in * in Tab. 6. This degradation highlights the importance to reduce visual tokens (Sec. 4) to adapt to multi-image scenarios.